Sarah invested $5000 in two funds, paying 7% and 9% annual interest respectively. If the total interest earned for the year was $375, how much was invested at each rate?
Found 2 solutions by stanbon, checkley77:
Answer by stanbon(75887)

Sarah invested $5000 in two funds, paying 7% and 9% annual interest respectively. If the total interest earned for the year was $375, how much was invested at each rate?
---------------------------
Interest Equation: 0.07x + 0.09(5000-x) = 375 dollars
7x + 9(5000-x) = 37500
7x + 45000 -9x = 37500
-2x = -7500
x = $3750.00 (amt. invested at 7%
5000-3750 = $1250 (amt. invested at 9%)
=========================================
Cheers,
Stan H.

Answer by checkley77(12844)

.09x+.07(5,000-x)=375
.09x+350-.07x=375
.02x=375-350
.02x=25
x=25/.02
x=$1,250 invested @ 9%.
5,000-1,250=$3,750 invested @ 7%.
Proof:
.09*1,250+.07*3,750=375
112.50+262.5=375
375=375